a stone is projected horizontally with 20 m per second from top of tall building calculate its position and velocity after 3 second neglecting the air resistance

Grade:11

1 Answers

Pawan joshi
5 years ago
Since horizontal component of velocity remains constant throughout the projectile but vertical component changes ,thus vertical velocity after 3s is Vy=Uy+at
Vy=10×3=30m/s. 
Net velocity, V=√Vx2+Vy2=√(20)2+(30)2=36m/s. 
Vertical displacement after 3s, y=1/2×10×3^2=45m
Horizontal displacement after 3s, x=20×3=60m